THE AB INITIO CALCULATION OF THE C_{3} − AR VAN DER WAALS COMPLEX

BON-GI LIN, SHIH-MING WEN, A. MEBEL, AND \underline{YEN-CHU HSU}, INSTITUTE OF ATOMIC AND MOLECULAR SCIENCES, ACADEMIA SINICA, P. O. BOX 23-166, TAIPEI 107, TAIWAN, R. O. C..

The equilibrium geometry of the ground electronic state of C_{3}Ar has been determined to be T-shaped at the level of CCSD(T)/cc-pVQZ: ℓ(C − C)=1.298Å, ^{}C̸ − C − C=173.75^{∘}, ℓ(C_{3} − Ar)=3.85Å, ^{}C̸_{3} − Ar=90^{∘}. Its binding energy is about 130 cm^{−1}and the vdW stretch is 30 cm^{−1}. The internal rotation of the complex is strongly coupled with the vdW stretch. Other conformers (two in-plane and one out-of-plane) have been identified as the C_{3} moiety of the complex bends away from the equlibrium angle. The computation results will be used to qualitatively interpret our experimental observations.